# Optimal Tests of the Composite Null Hypothesis Arising in Mediation Analysis

### v2

@@ Abstract
The indirect effect of an exposure on an outcome through an intermediate variable can be identified by a product of regression coefficients under certain causal and regression modeling assumptions. In this context, the null hypothesis of no indirect effect is a composite null hypothesis, as the null holds if either regression coefficient is zero. A consequence is that traditional hypothesis tests are severely underpowered near the origin (i.e., when both coefficients are small with respect to standard errors). We propose hypothesis tests that (i) preserve level alpha type 1 error, (ii) meaningfully improve power when both true underlying effects are small relative to sample size, and (iii) preserve power when at least one is not. One approach gives a closed-form test that is minimax optimal with respect to local power over the alternative parameter space. Another uses sparse linear programming to produce an approximately optimal test for a Bayes risk criterion. We discuss adaptations for performing large-scale hypothesis testing as well as modifications that yield improved interpretability. We provide an R package that implements the minimax optimal test.
@@ 1 INTRODUCTION
Mediation analysis is a widely popular discipline that seeks to understand the mechanism by which an exposure affects an outcome by learning about intermediate events that transmit part or all of its effect. For instance, if one is interested in the effect of an exposure $A$ on an outcome $Y$, one might posit that at least in part, $A$ affects $Y$ by first affecting some intermediate event $M$, which in turn affects $Y$. The effect along such a causal pathway is known as an indirect effect. Despite its long history and broad application, with an exception from recent work (van Garderen and van Giersbergen, 2022a, ), existing hypothesis tests of a single mediated effect are overly conservative and underpowered in a certain region of the alternative hypothesis space. In this article, we demonstrate the reason for this suboptimal behavior, and develop new hypothesis tests that deliver optimal power in some decision theoretic senses while preserving type 1 error uniformly over the composite null subspace of the parameter space. We focus on the case in which the indirect effect is identified by a product of coefficients or functions of coefficients that can be estimated with uniform joint convergence to a bivariate normal distribution with diagonal covariance matrix. In fact, our approach is not limited to the realm of mediation analysis, but can be applied to test for any product of two coefficients that are estimable in the above sense.
@@
Briefly, the inferential problem that arises in mediation analysis is that when the mediated effect is identified by a product of coefficients, even while the joint distribution of estimators of these coefficients may converge uniformly to a bivariate normal distribution, neither their product nor the minimum of their absolute values will converge uniformly to a Gaussian law. Thus, asymptotic approximation-based univariate test statistics that are a function of either of these summaries of the two coefficient estimates will have poor properties in an important region of the parameter space for any given finite sample.
Previously, MacKinnon et al., (2002) and Barfield et al., (2017) presented results from simulation studies comparing hypothesis tests of a mediated effect. Both articles observed that all existing tests are either underpowered in some scenarios or do not preserve nominal type 1 error in all scenarios. Both articles found the so-called joint significance test to be the best overall performing test, while Barfield et al., (2017) found a bootstrap-based test to have comparable performance across some scenarios, with the exception of the setting with a rare binary outcome. Huang, (2019) proved that the joint significance test yields a smaller $p$-value than those of both normality-based tests and normal product-based tests.
@@
Barfield et al., (2017) highlighted the problem of large-scale hypothesis testing for mediated effects, which has become a popular inferential task in genomics, where genomic measures such as gene expression or DNA methylation are hypothesized to be potential mediators of the effect of genes or other exposures on various outcomes. In fact, the conservative behavior of mediation hypothesis tests are most evident in large-scale hypothesis tests, since one can then compare the distribution of $p$-values with the uniform distribution. Huang, (2019) showed that the joint significance test and delta method-based tests both generate $p$-values forming a distribution that stochastically dominates the uniform distribution. However, while this behavior is readily observable in large-scale hypothesis tests, the problem is no different for a single hypothesis test; it is merely less obvious.
@@
Recently, Huang, (2019), Dai et al., (2020), Liu et al., (2022), and Du et al., (2023) developed large-scale hypothesis testing procedures for mediated effects that account for the conservativeness of standard tests. The latter compared these methods as well as traditional methods in a large simulation study. The method proposed by Huang, (2019) is based on a distributional assumption about the two coefficients among all distributions being tested that satisfy the null hypothesis. By contrast, the methods proposed by Dai et al., (2020), Liu et al., (2022), and Du et al., (2023) involve estimating the proportions of three different components of the composite null hypothesis: when both coefficients are zero, when one is zero and the other is not, and vice versa. Dai et al., (2020) and Du et al., (2023) then relaxed the threshold of the joint significance test and delta method-based test, respectively, based on these estimated proportions. Liu et al., (2022) used the estimated proportions to construct a new test statistic based on $p$-values attained under each of these components of the null hypothesis. Our proposed methods resolve the conservativeness issue across the entire null hypothesis space even for a single test, hence they require neither assumptions about the distributions of the coefficients nor estimation of proportions of components of the null hypothesis space. The latter is important, because in finite samples, there is no single distribution of a test statistic under the sub-case of the null hypothesis in which one parameter is null and the other is not. The distribution of a test statistic in this sub-case can vary from being well-approximated by the distribution when one parameter is zero and the other is infinite to being well-approximated by the distribution when both parameters are zero (indeed, this sub-case contains parameter values arbitrarily close to (0,0)). Instead, we can simply extend our method to the large-scale hypothesis test setting while adjusting for multiple testing using standard Bonferroni or Benjamini–Hochberg corrections.
@@
van Garderen and van Giersbergen, 2022a is the most closely related work to this article. The authors develop a test for a mediated effect based on defining a rejection region in $\mathbb{R}^{2}$. In their original version (van Garderen and van Giersbergen,, 2020), they prove that a similar test does not exist within a certain class of tests. In this article, we show that a similar test does in fact exist in a slightly expanded class of tests, and that it is the unique such test in this class (in a more recent version (van Garderen and van Giersbergen, 2022a, ), the authors now recognize the existence of this test). They focus on developing an almost-similar test, i.e., they attempt to get as close to a similar test as possible while preserving type 1 error by optimizing a particular objective function involving a penalty for deviations from the nominal type 1 error level. By contrast, we consider two approaches for optimizing decision theoretic criteria that characterize power while preserving type 1 error. The solution to one of these has a closed form and is an exactly similar test. The other is an approximate Bayes risk optimal test inspired by the test of Rosenblum et al., (2014).
@@
The remainder of the article is organized as follows. In Section 2, we formalize the problem and explain the shortcomings of traditional tests. In Sections 3 and 4, we present the minimax optimal and Bayes risk optimal test, respectively. In Section 5, we discuss adaptations for large-scale mediation hypothesis testing. In Section 6, we discuss interpretability challenges with the proposed tests, and introduce modifications leading to improved interpretability. In Section 7, we present results from simulation studies. In Section 8, we apply our methodology to two data sets: one to test whether cognition mediates the effect of cognitive remediation therapy on selt-esteem in patients with schizophrenia, and another to test many mediation hypotheses of whether a host of DNA methylation CpG sites mediate the effect of smoking status on lung function. In Section 9, we conclude with a discussion.
@@ 2 PRELIMINARIES
We begin with the general problem statement, which we will then connect to mediation analysis. Suppose we have $n$ independent, identically distributed (i.i.d.) observations from a distribution $P_{\bm{\delta}}$ indexed by the parameter $\bm{\delta}$, which contains two scalar parameters $\delta_{x}$ and $\delta_{y}$, and we wish to test the null hypothesis $H_{0}:``\delta_{x}\delta_{y}=0"$ against its alternative $H_{1}:``\delta_{x}\delta_{y}\neq 0"$. Further, suppose we have a uniformly asymptotically normal and unbiased estimator (Robins and Ritov,, 1997) $(\hat{\delta}_{x},\hat{\delta}_{y})$ for $(\delta_{x},\delta_{y})$, i.e.,
@@
as $n\rightarrow\infty$ for all $(t_{x},t_{y})^{\top}\in\mathbb{R}^{2}$, where $\bm{\Sigma_{n}}$ is a diagonal matrix that is a consistent estimator of the asymptotic covariance matrix, and $\Phi_{2}$ is the cumulative distribution function of the bivariate Gaussian distribution with identity covariance. We wish to test $H_{0}$ against $H_{1}$ with at most $\alpha$ type 1 error for each $(\delta_{x},\delta_{y})$ satisfying $\delta_{x}\delta_{y}=0$ and maximizing power (in some sense) everywhere else. Clearly, $H_{0}$ is a very particular type of composite null hypothesis, which in the space spanned by $\delta_{x}$ and $\delta_{y}$ consists of the $x$ and $y$ axes.
The composite null hypothesis $H_{0}$ arises naturally in mediation analysis under certain modeling assumptions. Suppose we observe $n$ i.i.d. copies of $(\bm{C}^{\top},A,M,Y)^{\top}$, where $A$ is the exposure of interest, $Y$ is the outcome of interest, $M$ is a potential mediator that is temporally intermediate to $A$ and $Y$, and $\bm{C}$ is a vector of baseline covariates that we will assume throughout to be sufficient to control for various sorts of confounding needed for the indirect effect to be identified. The natural indirect effect (NIE) is the mediated effect of $A$ on $Y$ through $M$. We formally define the NIE and discuss its identification in the Supplementary Materials. If the linear models with main effect terms given by
@@
are correctly-specified, then the identification formula for the natural indirect effect reduces to $\beta_{1}\theta_{2}$. The product method estimator is the product of estimates of $\beta_{1}$ and $\theta_{2}$ obtained by fitting the above regression models. Under models (2) and (3) and standard regularity conditions, the two factors of the product method estimator will satisfy the uniform joint convergence statement in (1). In fact, (1) will hold for a more general class of models, though there are important limitations to this class. For instance, consider the outcome model with exposure-mediator interaction replacing model (3):
@@
Under models (2) and (4), the natural indirect effect comparing treatment levels $a^{\prime}$ and $a^{\prime\prime}$ is identified by $(\theta_{2}+\theta_{3}a^{\prime})\beta_{1}(a^{\prime}-a^{\prime\prime})$, hence the estimator $(\hat{\theta}_{2}+\hat{\theta}_{3}a^{\prime})\hat{\beta}_{1}(a^{\prime}-a^{\prime\prime})$, where $\hat{\theta}_{2}$, $\hat{\theta}_{3}$, and $\hat{\beta}_{1}$ are estimated by fitting the linear regression models (2) and (4), will also satisfy (1) under standard regularity conditions, with $\delta_{x}=\theta_{2}+\theta_{3}a^{\prime}$ and $\delta_{y}=\beta_{1}(a^{\prime}-a^{\prime\prime})$.
@@
The problem with traditional tests of the composite null $H_{0}$ against its alternative $H_{1}$ stems from the fact that while $(\hat{\delta}_{x},\hat{\delta}_{y})$ converges uniformly in $(\delta_{x},\delta_{y})$, certain scalar functions of these estimators, e.g., the product method estimator $\hat{\delta}_{x}\hat{\delta}_{y}$, do not. This can be seen by examining the delta method approach proposed by Sobel, (1982) (often referred to as the Sobel test). When $(\delta_{x},\delta_{y})\neq(0,0)$, the delta method yields $\sqrt{n}(\hat{\delta}_{x}\hat{\delta}_{y}-\delta_{x}\delta_{y})\rightsquigarrow\mathcal{N}(0,\delta_{y}^{2}\sigma_{x}^{2}+\delta_{x}^{2}\sigma_{y}^{2})$, where $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the diagonal elements of $\bm{\Sigma}$. Thus, when precisely one of $\delta_{x}$ or $\delta_{y}$ is zero, $Z_{n}^{\text{prod}}:=n^{1/2}\hat{\delta}_{x}\hat{\delta}_{y}/(\hat{\delta}_{y}^{2}s_{x}^{2}+\hat{\delta}_{x}^{2}s_{y}^{2})^{1/2}\rightsquigarrow\mathcal{N}(0,1)$, where $s^{2}_{x}$ and $s^{2}_{y}$ are consistent estimates of $\sigma^{2}_{x}$ and $\sigma^{2}_{y}$, respectively. However, as $(\delta_{x},\delta_{y})\rightarrow(0,0)$, this no longer holds, as the denominator is also converging to zero. Liu et al., (2022) showed that this instead converges to a centered normal distribution with variance 1/4. Thus, the convergence in distribution of the delta method test statistic $Z_{n}^{\text{prod}}$ is not uniform, because for any given sample size, no matter how large, there will be a region in the parameter space around $(0,0)$ where the standard normal approximation can be very poor. This is especially problematic, as the region around $(0,0)$ is precisely where we would expect the truth to often lie. As one moves closer to the origin in the parameter space, the true distribution of $Z_{n}^{\text{prod}}$ begins to more closely resemble $N(0,1/4)$. This is demonstrated by a Monte Carlo sampling approximation of density functions in Figure S1.a in the Supplementary Materials, where we fix $\delta_{y}=0$ and vary $\delta_{x}$ from 0 to 0.3. The delta method is using an approximation based on the standard normal distribution, when in reality, the test statistic’s true distribution is much more concentrated about zero, as can be seen in the figure. This explains the conservative behavior of the delta method-based test near the origin.
Hypothesis tests based on inverting bootstrap confidence intervals are also known to perform poorly near the origin of the parameter space. The reason is that bootstrap theory requires continuity in the asymptotic distribution over the parameter space. As we have seen, there is a singularity in the pointwise asymptotic distribution at the origin.
@@
Another popular test of the composite null $H_{0}$ against its alternative $H_{1}$ is known as the joint significance test (Cohen et al.,, 2013), which is an intersection-union test based on the simple logic that both $\delta_{x}$ and $\delta_{y}$ must be nonzero for $\delta_{x}\delta_{y}$ to be nonzero. The joint significance test with nominal level $\alpha$ amounts to testing the two null hypotheses $H_{0}^{x}:``\delta_{x}=0"$ and $H_{0}^{y}:``\delta_{y}=0"$ against their alternatives $H_{1}^{x}:``\delta_{x}\neq 0"$ and $H_{1}^{y}:``\delta_{y}\neq 0"$ separately using the Wald statistics $Z^{x}_{n}:=\sqrt{n}\hat{\delta}_{x}/s_{x}$ and $Z^{y}_{n}:=\sqrt{n}\hat{\delta}_{y}/s_{y}$, and rejecting $H_{0}$ for $H_{1}$ only if $H_{0}^{x}$ and $H_{0}^{y}$ are both rejected for their alternatives by the corresponding Wald tests with nominal level $\alpha$. As we previously noted, this test has been shown to perform better in simulations than other common tests of $H_{0}$ against $H_{1}$. However, it still suffers from the lack of power near the origin of the parameter space we see in other tests. This is once again due to a lack of uniform convergence near the origin. The joint significance test implicitly relies on the approximation $Z_{n}^{\text{min}}:=|Z_{n}^{x}|\wedge|Z_{n}^{y}|\sim|\mathcal{N}|(0,1)$, where $|\mathcal{N}|(0,1)$ represents the folded normal distribution with mean zero and variance one. However, this approximation is poor when $\delta_{x}$ and $\delta_{y}$ are both close to zero, as is shown in Figure S1.b in the Supplementary Materials, where we once again fix $\delta_{y}=0$ and vary $\delta_{x}$ from 0 to 0.3.
In order to avoid the problems that arise due to non-uniform convergence, we will instead focus on constructing a rejection region in the parameter space spanned by the vector $(\delta_{x},\delta_{y})$ in $\mathbb{R}^{2}$, rather than the parameter space spanned by the product $\delta_{x}\delta_{y}$ in $\mathbb{R}$. In the former, we know convergence is uniform and there are no discontinuities in the asymptotic distribution over the parameter space. Studying the rejection regions of the traditional tests in $\mathbb{R}^{2}$ illustrates their conservative behavior from another perspective.
@@
Consider a sequence of local noncentrality parameters with respect to a given direction $(\delta_{x},\delta_{y})^{\top}$, $(\delta_{x,n},\delta_{y,n})^{\top}:=(\delta_{x}/n^{1/2},\delta_{y}/n^{1/2})^{\top}$, and define $(\delta_{x}^{*},\allowbreak\delta_{y}^{*})^{\top}\allowbreak:=\allowbreak\mathrm{plim}_{n\rightarrow\infty}\allowbreak n^{1/2}\allowbreak\bm{\Sigma_{n}}^{-1/2}\allowbreak(\delta_{x,n},\allowbreak\delta_{y,n})^{\top}\allowbreak=\allowbreak\bm{\Sigma}^{-1/2}\allowbreak(\delta_{x},\allowbreak\delta_{y})^{\top}$ where $\mathrm{plim}$ stands for the limit in probability. Further, define the bivariate test statistic $(Z^{x}_{n},Z^{y}_{n})^{\top}:=n^{1/2}\bm{\Sigma_{n}}^{-1/2}(\hat{\delta}_{x},\hat{\delta}_{y})^{\top}$ so that $(Z^{x}_{n},Z^{y}_{n})^{\top}$ is approximately distributed as $\mathcal{N}\{(\delta_{x}^{*},\allowbreak\delta_{y}^{*})^{\top},\allowbreak\bm{I_{2}}\}$. Let $(Z_{*}^{x},Z_{*}^{y})^{\top}$ be a random vector following this limiting distribution exactly. Based on this asymptotic approximation, the probability of the Wald test rejecting $H_{0}^{x}$ for $H_{1}^{x}$ at nominal level $\alpha$ is $\alpha$ under all elements of $\{(\delta_{x}^{*},\delta_{y}^{*}):\delta_{x}^{*}=0\}$, and likewise the probability of the Wald test rejecting $H_{0}^{y}$ for $H_{1}^{y}$ at nominal level $\alpha$ is $\alpha$ under all elements of $\{(\delta_{x}^{*},\delta_{y}^{*}):\delta_{x}^{*}=0\}$. The rejection region in $\mathbb{R}^{2}$ corresponding to the joint significance test is the intersection of these two rejection regions, as shown in the plot in Figure 1(a).

### v3

@@ Abstract
The indirect effect of an exposure on an outcome through an intermediate variable can be identified by a product of two regression coefficients under certain causal and regression modeling assumptions. In this context, the null hypothesis of no indirect effect is a composite null hypothesis, as the null holds if either regression coefficient is zero. A consequence is that traditional hypothesis tests are severely underpowered near the origin (i.e., when both coefficients are small with respect to standard errors). We propose hypothesis tests that (i) preserve level alpha type 1 error, (ii) meaningfully improve power when both true underlying effects are small relative to sample size, and (iii) preserve power when at least one is not. One approach gives a closed-form test that is minimax optimal with respect to local power over the alternative parameter space. Another uses sparse linear programming to produce an approximately optimal test for a Bayes risk criterion. We discuss adaptations for performing large-scale hypothesis testing as well as modifications that yield improved interpretability. We provide an R package that implements our proposed methodology.
@@ 1 INTRODUCTION
Mediation analysis is a widely popular discipline that seeks to understand the mechanism by which an exposure affects an outcome by learning about intermediate events that transmit part or all of its effect. For instance, if one is interested in the effect of an exposure $A$ on an outcome $Y$, one might posit that at least in part, $A$ affects $Y$ by first affecting some intermediate event $M$, which in turn affects $Y$. Despite its long history and broad application, with an exception from recent work (van Garderen and van Giersbergen,, 2024), existing hypothesis tests of a single mediated effect are overly conservative and underpowered in a certain region of the alternative hypothesis space. In this article, we demonstrate the reason for this suboptimal behavior, and develop new hypothesis tests that deliver optimal power in some decision theoretic senses while preserving type 1 error uniformly over the composite null subspace of the parameter space. We focus on the case in which the indirect effect is identified by a product of two coefficients or functions of coefficients that can be estimated with uniform joint convergence to a bivariate normal distribution with diagonal covariance matrix. In fact, our approach is not limited to the realm of mediation analysis, but can be applied to test for any product of two coefficients that are estimable in the above sense.
@@
Briefly, the inferential problem that arises in mediation analysis is that when the mediated effect is identified by a product of two coefficients, even while the joint distribution of estimators of these coefficients may converge uniformly to a bivariate normal distribution, neither their product nor the minimum of their absolute values will converge uniformly to a Gaussian law. Thus, asymptotic approximation-based univariate test statistics that are a function of either of these summaries of the two coefficient estimates will have poor properties in an important region of the parameter space for any given finite sample.
@@
Previously, MacKinnon et al., (2002) and Barfield et al., (2017) presented simulation studies comparing hypothesis tests of a mediated effect. Both articles observed that all existing tests are either underpowered in some scenarios or do not preserve nominal type 1 error in all scenarios. Both articles found the so-called joint significance test to be the best overall performing test, while Barfield et al., (2017) found a bootstrap-based test to have comparable performance across some scenarios. Huang, (2019) proved that the joint significance test yields a smaller $p$-value than those of both normality-based tests and normal product-based tests. Barfield et al., (2017) highlighted the problem of large-scale hypothesis testing for mediated effects, which has become a popular inferential task in genomics, where genomic measures such as gene expression or DNA methylation are hypothesized to be potential mediators of the effect of genes or other exposures on various outcomes. In fact, the conservative behavior of mediation hypothesis tests is most evident in large-scale hypothesis testing, since one can compare the distribution of $p$-values with the uniform distribution. Huang, (2019) showed that the joint significance and delta method tests both generate $p$-values forming a distribution that stochastically dominates the uniform distribution. However, while this behavior is readily observable in large-scale hypothesis testing, the problem is no different for a single hypothesis test; it is merely less obvious.
@@
Recently, Huang, (2019), Dai et al., (2022), Liu et al., (2022), and Du et al., (2023) developed large-scale hypothesis testing procedures for mediated effects that account for the conservativeness of standard tests. The latter compared these methods as well as traditional methods in a large simulation study. The method proposed by Huang, (2019) is based on a distributional assumption about the two coefficients among all distributions being tested that satisfy the null hypothesis. The methods of Dai et al., (2022), Liu et al., (2022), and Du et al., (2023) involve estimating the proportions of three components of the composite null hypothesis: when both coefficients are zero, when one is zero and the other is not, and vice versa. Dai et al., (2022) and Du et al., (2023) then relaxed the threshold of the joint significance test and delta method-based test, respectively, based on these estimated proportions. Liu et al., (2022) used the estimated proportions to construct a new test statistic based on $p$-values attained under each of these components of the null hypothesis. Our proposed methods resolve the conservativeness issue across the entire null hypothesis space even for a single test, hence they require neither assumptions about the distributions of the coefficients nor estimation of proportions of components of the null hypothesis space. The latter is important, because in finite samples, there is no single distribution of a test statistic under the sub-case of the null hypothesis in which one parameter is null and the other is not. The distribution of a test statistic in this sub-case can vary from being well-approximated by the distribution when one parameter is zero and the other is infinite to being well-approximated by the distribution when both parameters are zero. Instead, we can simply extend our method to the large-scale hypothesis test setting while adjusting for multiple testing using standard Bonferroni or Benjamini–Hochberg corrections.
@@
van Garderen and van Giersbergen, (2024) is the most closely related work to this article. The authors develop a test for a mediated effect based on defining a rejection region in $\mathbb{R}^{2}$. In their original version (van Garderen and van Giersbergen,, 2020), they proved that a similar test does not exist within a certain class of tests. In this article, we show that a similar test does in fact exist in a slightly expanded class of tests, and that it is the unique such test in this class (in the more recent version, the authors now recognize the existence of this test). They focus on developing an almost-similar test, i.e., they approximate a similar test by optimizing a particular objective function involving a penalty for deviations from the nominal type 1 error level. By contrast, we consider two approaches for optimizing decision theoretic criteria that characterize power while preserving type 1 error. The solution to one of these has a closed form and is an exactly similar test. The other is an approximate Bayes risk optimal test inspired by the test of Rosenblum et al., (2014).
@@
The remainder of the article is organized as follows. In Section 2, we formalize the problem and explain the shortcomings of traditional tests. In Sections 3 and 4, we present the minimax optimal and Bayes risk optimal tests, respectively. In Section 5, we discuss adaptations for large-scale mediation hypothesis testing. In Section 6, we discuss interpretability challenges with the proposed tests, and introduce modifications leading to improved interpretability. In Section 7, we present results from simulation studies. In Section 8, we apply our methodology to two data sets: one to test whether cognition mediates the effect of cognitive remediation therapy on self-esteem in patients with schizophrenia, and another to test many mediation hypotheses of whether a host of DNA methylation CpG sites mediate the effect of smoking status on lung function. In Section 9, we conclude with a discussion.
@@ 2 PRELIMINARIES
We begin with the general problem statement, which we will then connect to mediation analysis. Suppose we have $n$ independent, identically distributed (i.i.d.) observations from a distribution $P_{\bm{\delta}}$ indexed by the parameter $\bm{\delta}=(\delta_{x},\delta_{y})^{\top}\in\mathbb{R}^{2}$, and we wish to test the null hypothesis $H_{0}:``\delta_{x}\delta_{y}=0"$ against its alternative $H_{1}:``\delta_{x}\delta_{y}\neq 0"$. Further, suppose we have a uniformly asymptotically normal and unbiased estimator (Robins and Ritov,, 1997) $(\hat{\delta}_{x},\hat{\delta}_{y})$ for $(\delta_{x},\delta_{y})$, i.e.,
@@
as $n\rightarrow\infty$ for all $(t_{x},t_{y})^{\top}\in\mathbb{R}^{2}$, where $\bm{\Sigma_{n}}$ is a diagonal matrix that is a consistent estimator of the asymptotic covariance matrix of $(\hat{\delta}_{x},\hat{\delta}_{y})^{\top}$, and $\Phi_{2}$ is the cumulative distribution function of the bivariate Gaussian distribution with identity covariance. We wish to test $H_{0}$ against $H_{1}$ with at most $\alpha$ type 1 error for each $(\delta_{x},\delta_{y})$ satisfying $\delta_{x}\delta_{y}=0$ and maximizing power elsewhere. Clearly, $H_{0}$ is a very particular type of composite null hypothesis, which in the space spanned by $\delta_{x}$ and $\delta_{y}$ consists of the $x$ and $y$ axes.
The composite null hypothesis $H_{0}$ arises naturally in mediation analysis under certain modeling assumptions. Suppose we observe $n$ i.i.d. copies of $(\bm{C}^{\top},A,M,Y)^{\top}$, where $A$ is the exposure of interest, $Y$ is the outcome of interest, $M$ is a potential mediator, and $\bm{C}$ is a vector of baseline covariates that is assumed to control for confounding such that the indirect effect is identified. The natural indirect effect (NIE) is the mediated effect of $A$ on $Y$ through $M$. We formally define the NIE and discuss its identification in the Supplementary Materials. If the linear models with main effect terms given by
@@
are correctly-specified, then the identification formula for the natural indirect effect reduces to $\beta_{1}\theta_{2}$. The product method estimator is the product of estimates of $\beta_{1}$ and $\theta_{2}$ obtained by fitting the above regression models. Under models (2) and (3) and standard regularity conditions, the two factors of the product method estimator will jointly satisfy the uniform convergence statement in (1). In fact, (1) will hold for a more general class of models, though there are important limitations to this class. For instance, consider the outcome model with exposure-mediator interaction replacing model (3):
@@
Under models (2) and (4), the natural indirect effect comparing treatment levels $a^{\prime}$ and $a^{\prime\prime}$ is identified by $(\theta_{2}+\theta_{3}a^{\prime})\beta_{1}(a^{\prime}-a^{\prime\prime})$, which is naturally estimated by $(\hat{\theta}_{2}+\hat{\theta}_{3}a^{\prime})\hat{\beta}_{1}(a^{\prime}-a^{\prime\prime})$, where $\hat{\theta}_{2}$, $\hat{\theta}_{3}$, and $\hat{\beta}_{1}$ are estimated by fitting the linear regression models (2) and (4). Setting $\delta_{x}=\theta_{2}+\theta_{3}a^{\prime}$, $\delta_{y}=\beta_{1}(a^{\prime}-a^{\prime\prime})$, $\hat{\delta}_{x}=\hat{\theta}_{2}+\hat{\theta}_{3}a^{\prime}$, and $\hat{\delta}_{y}=\hat{\beta}_{1}(a^{\prime}-a^{\prime\prime})$, the joint estimator will also satisfy (1) under standard regularity conditions.
@@
The problem with traditional tests of the composite null $H_{0}$ against its alternative $H_{1}$ stems from the fact that while $(\hat{\delta}_{x},\hat{\delta}_{y})$ converges uniformly in $\bm{\delta}$, certain scalar functions of these estimators, e.g., the product method estimator $\hat{\delta}_{x}\hat{\delta}_{y}$, do not. This can be seen by examining the delta method approach proposed by Sobel, (1982) (often referred to as the Sobel test). When $\bm{\delta}\neq(0,0)^{\top}$, the delta method yields $\sqrt{n}(\hat{\delta}_{x}\hat{\delta}_{y}-\delta_{x}\delta_{y})\rightsquigarrow\mathcal{N}(0,\delta_{y}^{2}\sigma_{x}^{2}+\delta_{x}^{2}\sigma_{y}^{2})$, where $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the diagonal elements of the limit in probability of $\bm{\Sigma}_{n}$. Thus, when precisely one of $\delta_{x}$ or $\delta_{y}$ is zero, $Z_{n}^{\text{prod}}:=n^{1/2}\hat{\delta}_{x}\hat{\delta}_{y}/(\hat{\delta}_{y}^{2}s_{x}^{2}+\hat{\delta}_{x}^{2}s_{y}^{2})^{1/2}\rightsquigarrow\mathcal{N}(0,1)$, where $s^{2}_{x}$ and $s^{2}_{y}$ are consistent estimates of $\sigma^{2}_{x}$ and $\sigma^{2}_{y}$, respectively. However, as $\bm{\delta}\rightarrow(0,0)^{\top}$, this no longer holds, as the denominator also goes to zero. Liu et al., (2022) showed that this instead converges to a centered normal distribution with variance 1/4. Thus, the convergence in distribution of the delta method test statistic $Z_{n}^{\text{prod}}$ is not uniform, and for any given sample size, there will be a region in the parameter space around $(0,0)$ where the standard normal approximation is poor. This is especially problematic, as the region around $(0,0)$ is precisely where we expect the truth to often lie. As $\bm{\delta}$ approaches $(0,0)$, $Z_{n}^{\text{prod}}$ begins to more closely resemble $N(0,1/4)$, as we show in a Monte Carlo sampling approximation of density functions in Figure S1.a in the Supplementary Materials, where we fix $\delta_{y}=0$ and vary $\delta_{x}$ from 0 to 0.3. The delta method uses an approximation based on the standard normal distribution, when in reality, the test statistic’s true distribution is much more concentrated about zero, as shown in the figure. Hypothesis tests based on inverting bootstrap confidence intervals are also known to perform poorly near the origin of the parameter space due to the singularity in the pointwise asymptotic distribution at the origin.
@@
Another popular test of the composite null $H_{0}$ against its alternative $H_{1}$ is known as the joint significance test (Cohen et al.,, 2013), which is based on the logic that both $\delta_{x}$ and $\delta_{y}$ must be nonzero for $\delta_{x}\delta_{y}$ to be nonzero. The joint significance test with nominal level $\alpha$ amounts to testing the two null hypotheses $H_{0}^{x}:``\delta_{x}=0"$ and $H_{0}^{y}:``\delta_{y}=0"$ against their alternatives $H_{1}^{x}:``\delta_{x}\neq 0"$ and $H_{1}^{y}:``\delta_{y}\neq 0"$ separately using the Wald statistics $Z^{x}_{n}:=\sqrt{n}\hat{\delta}_{x}/s_{x}$ and $Z^{y}_{n}:=\sqrt{n}\hat{\delta}_{y}/s_{y}$, and rejecting $H_{0}$ for $H_{1}$ only if $H_{0}^{x}$ and $H_{0}^{y}$ are both rejected for their alternatives with nominal level $\alpha$. As noted, this test has been shown to perform better in simulations than other common tests of $H_{0}$ against $H_{1}$. However, it still suffers from the lack of power near the origin of the parameter space we see in other tests. This is again due to a lack of uniform convergence near the origin. The joint significance test implicitly relies on the approximation $Z_{n}^{\text{min}}:=|Z_{n}^{x}|\wedge|Z_{n}^{y}|\sim|\mathcal{N}|(0,1)$, where $|\mathcal{N}|(0,1)$ represents the folded standard normal distribution. This approximation is poor when $\delta_{x}$ and $\delta_{y}$ are both close to zero, as is shown in Figure S1.b in the Supplementary Materials.
@@
To avoid the problems that arise due to non-uniform convergence, we instead construct a rejection region in the parameter space spanned by the vector $(\delta_{x},\delta_{y})$ in $\mathbb{R}^{2}$, rather than that spanned by $\delta_{x}\delta_{y}$ in $\mathbb{R}$. In the former, convergence is uniform and there are no discontinuities in the asymptotic distribution. Studying the rejection regions of the traditional tests in $\mathbb{R}^{2}$ illustrates their conservative behavior from another perspective. Consider a sequence of local noncentrality parameters with respect to a given direction $(\delta_{x},\delta_{y})^{\top}$, $(\delta_{x,n},\delta_{y,n})^{\top}:=(\delta_{x}/n^{1/2},\delta_{y}/n^{1/2})^{\top}$, and define $\bm{\delta}^{*}=(\delta_{x}^{*},\allowbreak\delta_{y}^{*})^{\top}\allowbreak:=\allowbreak\mathrm{plim}_{n\rightarrow\infty}\allowbreak n^{1/2}\allowbreak\bm{\Sigma_{n}}^{-1/2}\allowbreak(\delta_{x,n},\allowbreak\delta_{y,n})^{\top}\allowbreak=\allowbreak\bm{\Sigma}^{-1/2}\allowbreak(\delta_{x},\allowbreak\delta_{y})^{\top}$ where $\mathrm{plim}$ is the limit in probability and $\bm{\Sigma}:=\mathrm{plim}_{n\rightarrow\infty}\bm{\Sigma}_{n}$. Define the bivariate test statistic $(Z^{x}_{n},Z^{y}_{n})^{\top}:=n^{1/2}\bm{\Sigma_{n}}^{-1/2}(\hat{\delta}_{x},\hat{\delta}_{y})^{\top}$ so that $(Z^{x}_{n},Z^{y}_{n})^{\top}$ is approximately distributed as $\mathcal{N}\{(\delta_{x}^{*},\allowbreak\delta_{y}^{*})^{\top},\allowbreak\bm{I_{2}}\}$, and let $(Z_{*}^{x},Z_{*}^{y})^{\top}\sim\mathcal{N}\{(\delta_{x}^{*},\allowbreak\delta_{y}^{*})^{\top},\allowbreak\bm{I_{2}}\}$. Based on this asymptotic approximation, the probability of the Wald test rejecting $H_{0}^{x}$ for $H_{1}^{x}$ at nominal level $\alpha$ is $\alpha$ under all elements of $\{\bm{\delta}^{*}:\delta_{x}^{*}=0\}$, and likewise for the Wald test rejecting $H_{0}^{y}$ for $H_{1}^{y}$ for all $\{\bm{\delta}^{*}:\delta_{y}^{*}=0\}$. The rejection region in $\mathbb{R}^{2}$ corresponding to the joint significance test is the intersection of these two rejection regions, as shown in the plot in Figure 1(a).